OpenStudy (anonymous):

Put the numbers in order, then find the median: -1/2, 0.81, 8.9, 1.5, 1.75, 1.32 A. 1.5 B. 1.41 C. 8.9

OpenStudy (anonymous):

you would take the two middle numbers 1.32 and 1.5 add them together then divide by 2. 1.32+1.5=2.82/2=1.41 so the median is 1.41